# SDFLoRA: Selective Dual-Module LoRA for Federated Fine-tuning with Heterogeneous Clients

### v1

@@ 3.4 Privacy-Aware Optimization (DP Compatibility)
where $C$ is the clipping norm and $\sigma$ is the noise multiplier. The perturbed gradients update $\theta_{k}^{(g)}$, which is then uploaded to the server. The local module $\theta_{k}^{(l)}$ is updated without noise, preserving client-specific expressiveness and improving optimization stability.
@@
Injecting noise into the entire adapter amplifies variance and can destabilize learning, especially for heterogeneous clients. By restricting DP to the global module, our framework protects transferable knowledge while avoiding unnecessary corruption of client-specific adaptations, yielding a better utility–privacy trade-off.
@@ 3.5 Algorithm Summary
Algorithm 1 summarizes one communication round of our method. Each client receives the global module, performs local training with dual-module adapters (optionally with DP on the global module), uploads only the global-module parameters, and the server performs selective stacking aggregation to produce the next global module.
Algorithm 1 Dual-Module Federated LoRA with Selective Stacking (one round)
Input: Frozen backbone $W$; global module $\theta^{(g,t)}$; client set $S_{t}$; weights $\{p_{k}\}$; rank budget $r_{\max}$ ; DP params $(C,\sigma)$
Output: Updated global module $\theta^{(g,t+1)}$
1: Server broadcasts $\theta^{(g,t)}$ to clients $k\in S_{t}$.
2:for each client $k\in S_{t}$ in parallel do
3:  Initialize dual-module adapters with $\theta^{(g,t)}$ and local state $\theta_{k}^{(l,t)}$.
4:  for local step $e=1,\dots,E$ do
5:   Compute gradients on batch; update $\theta_{k}^{(g)}$ (with DP-SGD if enabled) and $\theta_{k}^{(l)}$.
6:  end for
7:  Upload $\theta_{k}^{(g,t)}$ to server; keep $\theta_{k}^{(l,t)}$ locally.
@@
9: Server aggregates global modules via selective stacking: $\theta^{(g)}\leftarrow\mathrm{StackAgg}(\{\theta_{k}^{(g,t)}\}_{k\in S_{t}},\{p_{k}\})$.
10: Re-compress to rank budget $r_{\max}$ to obtain $\theta^{(g,t+1)}$; otherwise set $\theta^{(g,t+1)}\leftarrow\theta^{(g)}$.
@@ 4.1 Experimental Setup and Baselines
We evaluate federated parameter-efficient fine-tuning on GLUE-style NLU benchmarks Wang et al. (2018). We report results on QNLI, RTE, QQP, MNLI, and SST-2, which cover natural language inference, paraphrase identification, and sentiment classification. We use the official train/dev splits and report dev-set performance.
@@
We adopt LLaMA-7B as the frozen backbone and insert LoRA adapters Hu et al. (2022) into linear projection layers. Only LoRA parameters are optimized while the backbone weights remain fixed.
@@
We simulate a cross-device FL setting with $K=8$ clients and run $T=30$ communication rounds. Unless otherwise stated, all clients participate in each round and the server aggregates client updates with weights proportional to local data sizes, i.e., $\alpha_{k}=n_{k}/\sum_{j}n_{j}$. To model structural heterogeneity induced by rank mismatch, we set client ranks as $\{r_{k}\}=[4,4,8,8,8,8,16,16]$. When a homogeneous setting is required (e.g., for FedAvg-style baselines), we use a shared rank $r_{k}=r$ for all clients. In an additional scalability study, we vary the number of participating clients $K$ while keeping the remaining hyperparameters consistent.
@@
Each client uses Adam with a fixed learning rate of $2\times 10^{-4}$, batch size $128$, and trains for $200$ epochs. Within each communication round, clients perform $10$ local update steps before uploading their LoRA updates. All remaining optimizer hyperparameters follow the default settings in our implementation. These hyperparameters are kept identical across methods unless the experiment explicitly controls a specific factor (e.g., aggregation strategy or privacy budget), ensuring fair comparisons. [^1] The above hyperparameters follow the experimental configuration summarized in Table 1.
@@
We compare the proposed aggregation strategy against representative alternatives under identical training budgets: (i) Zero-padding aligns heterogeneous LoRA updates by padding lower-rank factors to the maximum rank before weighted averaging; (ii) FedAvg-style averaging performs standard weighted averaging of LoRA parameters in the homogeneous-rank setting; (iii) Stacking concatenates client low-rank updates along the rank dimension to preserve the union of client subspaces.
@@
Item Setting
Backbone LLaMA-7B
Adapter LoRA
Tasks QNLI, RTE, QQP, MNLI, SST-2
Clients ($K$) 8
Rounds ($T$) 30
Rank (hetero.) $\{r_{k}\}=[4,4,8,8,8,8,16,16]$
Aggregation $\alpha_{k}=n_{k}/\sum_{j}n_{j}$
Optimizer Adam
Learning rate $2\times 10^{-4}$
Batch size 128
Local epochs 200
Local steps/round 10
[…]
Software Ubuntu 22.04.5, Python 3.10, CUDA 12.4
Table 1: Experimental configuration and system environment.
@@ 4.2 Results under Rank and Data Heterogeneity
Rank-heterogeneous clients.
Table 2 compares stacking and zero-padding under mismatched client ranks. Stacking yields consistent and often substantial improvements across all four tasks. In particular, it improves QNLI from 95.81 to 96.96 (1.15%), QQP from 86.72 to 87.97 (1.25%), and shows especially large gains on RTE (4.26%) and MNLI (6.71%). The much larger gains on RTE and MNLI suggest that these tasks are more sensitive to directional misalignment across clients, where important low-rank update directions from some clients are severely distorted by zero-padding.
This behavior can be explained by how the two aggregation strategies treat heterogeneous ranks. Zero-padding forces all client updates into a common high-rank space by inserting artificial zero directions for low-rank clients, which effectively dilutes their informative components during weighted averaging. In contrast, stacking concatenates client factors along the rank dimension, preserving each client’s effective update subspace and preventing destructive interference between heterogeneous low-rank directions. As a result, stacking produces a more faithful reconstruction of the global update under rank mismatch.
Hetero Homo
Task Ours Padding Ours FedAvg Ours FedAvg
Rank – Rank = 4 Rank = 8
QNLI 96.96 95.81 97.78 95.81 98.47 96.50
RTE 99.71 95.45 88.89 85.94 90.05 85.62
QQP 87.97 86.72 87.60 84.40 85.30 83.10
MNLI(m) 72.19 65.48 73.05 70.20 73.60 69.85
MNLI($m$) 71.62 64.90 72.48 69.55 73.05 69.20
SST-2 82.30 81.10 82.14 82.06 81.88 81.90
Table 2: Fed aggregation across rank-heterogeneous and rank-homogeneous clients.
Homogeneous-rank sanity check.
To verify that the benefit of stacking is not merely due to resolving dimensional incompatibility, we further compare it with FedAvg-style averaging when all clients use the same rank (Table 2). Even in this homogeneous setting, stacking remains consistently superior or comparable.
For rank $r=4$, stacking improves QQP by 3.20%, QNLI by 1.97% , and RTE by 2.95% , while maintaining comparable performance on SST-2. For rank $r=8$, stacking again improves QQP (2.20%), QNLI (1.97%), and RTE (4.43%), with only a negligible difference on SST-2 (-0.02%). These results indicate that stacking acts as a subspace-preserving aggregator that is beneficial even when rank is fixed, rather than being a mere workaround for heterogeneous dimensions.
@@
We further study how the federated population size affects performance by varying the number of participating clients $K$ while keeping the training settings and optimization hyperparameters consistent. Figure 2 shows a clear scale benefit when moving from small to medium client populations: with too few clients (e.g., K=4K{=}4), global aggregation is less effective, which is consistent with the limited coverage of data distributions and update directions in small federations. As $K$ increases, accuracy improves substantially, indicating that a larger client pool provides richer and more diverse signals for the global model to generalize.
In particular, the curve is not strictly monotonic; after initial gains, we observe a stage-wise fluctuation (e.g., a temporary drop around K=16K{=}16 for one rank setting) before recovering at larger $K$. This behavior is expected in non-IID federated optimization, where enlarging the population simultaneously increases data coverage and amplifies cross-client update discrepancy, which may transiently destabilize aggregation. Comparing the two rank configurations in Figure 2, a lower rank exhibits better robustness in the small-to-medium regime, whereas a higher rank tends to achieve a higher ceiling once the client population becomes sufficiently large. Overall, this scalability study complements Table LABEL:tab:main_heteroLABEL:tab:dirichlet_std by showing that our method remains effective when the federation size varies and that structural capacity (rank) interacts non-trivially with the client population.
@@
We further examine robustness under Non-IID data partitions controlled by a Dirichlet parameter $\alpha$ using the client-level accuracy standard deviation in Table LABEL:tab:dirichlet_std. Lower variance indicates more consistent performance across heterogeneous clients.
@@
At $\alpha=1.0$, stacking remains slightly more stable on RTE (0.088 vs. 0.092), but exhibits higher variance on SST-2 and QNLI. This suggests that when data heterogeneity is moderate, preserving a broader set of client-specific low-rank directions may amplify inter-client discrepancies on certain tasks, even though the global model remains competitive. Importantly, when the partition approaches near-IID ($\alpha=10$), stacking again achieves uniformly lower variance across all tasks, e.g., reducing QNLI variance from 0.045 to 0.014, indicating that stacking does not inherently harm stability.
Figure 3: Client-level performance comparison between single-module and dual-module adapters under the same training budget. The dual-module design consistently improves accuracy across clients, validating the effectiveness of separating global and local low-rank adaptations.
@@ 4.3 Privacy and Rank Sensitivity
Privacy–utility trade-off.
We apply DP-SGD during local training and report task accuracies under different privacy budgets $\epsilon$ (Figure 4). As $\epsilon$ increases (weaker privacy), the required noise scale decreases and performance improves.
Figure 4: Accuracy under different privacy budgets $\epsilon$. Fixed denotes applying structural constraints (selective stacking and rank control) on the global module, while Unfixed corresponds to unconstrained aggregation over all updates. The fixed design exhibits substantially better stability under strong privacy constraints.
Figure 4 further shows that enforcing structural constraints on the global module significantly improves robustness under strong privacy constraints. When $\epsilon$ is small, unconstrained aggregation suffers from severe performance degradation, whereas the fixed global-module design maintains substantially higher accuracy.
@@
Tables 3 analyze how the LoRA rank affects performance. Without DP, smaller ranks can be more stable on some tasks, while overly large ranks may introduce redundancy and hurt aggregation. With DP ($\epsilon=1$), the optimal rank can shift due to noise, but small-to-medium ranks often remain a good trade-off.
@@
4.4 Evidence for Dual-Module Benefit
Finally, we include direct evidence that introducing a dual-adapter (dual-module) design improves federated fine-tuning. Under matched training budgets, the dual-adapter achieves a higher average performance than a single-adapter baseline, indicating that dual modules better exploit complementary low-rank subspaces.
@@ 5 Conclusion
We studied federated fine-tuning of large language models with LoRA under practical client heterogeneity, where clients may differ not only in data and systems but also in the rank configurations of low-rank adapters and privacy requirements. While stacking-based aggregation provides a natural way to reconcile rank mismatch, we argued that stacking alone is insufficient because it implicitly forces global alignment on all client updates, potentially suppressing personalization and becoming fragile under differential privacy noise.
To address these limitations, we proposed SDFLoRA framework that structurally decomposes each client adapter into a global module capturing transferable knowledge and a local module preserving client-specific adaptations. Building on this decomposition, we introduced selective stacking that applies stacking-based alignment only to the global module while keeping local modules private and unaggregated. This structural restriction mitigates over-alignment, improves robustness under rank heterogeneity, and naturally supports privacy-aware optimization by injecting DP noise exclusively into the global module.
Experiments on GLUE benchmarks under heterogeneous-rank settings demonstrated that our approach consistently outperforms representative federated LoRA baselines and achieves a better utility–privacy trade-off under DP constraints. In future work, we plan to extend the proposed structural separation to other adapter families and investigate adaptive mechanisms for automatically allocating global versus local capacity based on client characteristics and task demands.

### v2

@@ 3.4 Privacy Aware Optimization (DP Compatibility)
where $C$ is the clipping norm and $\sigma$ is the noise multiplier. The perturbed gradients update $\theta_{k}^{(g)}$, which is then uploaded to the server. The private module $\theta_{k}^{(l)}$ is updated without noise, preserving client-specific expressiveness and improving optimization stability.
@@
Injecting noise into the entire adapter amplifies variance and can destabilize learning, especially for heterogeneous clients. By restricting DP to the shared module, our framework protects transferable knowledge while avoiding unnecessary corruption of client-specific adaptations, yielding a better balance of utility and privacy.
@@ 3.5 Algorithm Summary
Algorithm 1 summarizes one communication round of our method. Each client receives the shared module, performs local training with decoupled adapters, uploads only the shared-module parameters, and the server performs selective stacking aggregation to produce the next shared module.
Algorithm 1 Selective Decoupled Federated LoRA (One Round)
1:Input: Frozen backbone $W$; shared module $\theta^{(g,t)}$; client set $S_{t}$; weights $\{p_{k}\}$; rank budget $r_{\max}$; DP parameters $(C,\sigma)$
2:Output: Updated shared module $\theta^{(g,t+1)}$
3: Server broadcasts $\theta^{(g,t)}$ to clients $k\in S_{t}$.
4:for each client $k\in S_{t}$ in parallel do
5:  Initialize decoupled adapters with $\theta^{(g,t)}$ and local state $\theta_{k}^{(l,t)}$.
6:  for local step $e=1,\dots,E$ do
7:   Compute gradients on batch; update $\theta_{k}^{(g)}$ (with DP-SGD if enabled) and $\theta_{k}^{(l)}$.
@@
9:  Upload $\theta_{k}^{(g,t)}$ to server; keep $\theta_{k}^{(l,t)}$ locally.
10:end for
11: Server aggregates shared modules via selective stacking:
12:   $\theta^{(g)}\leftarrow\mathrm{StackAgg}(\{\theta_{k}^{(g,t)}\}_{k\in S_{t}},\{p_{k}\})$.
13: Re-compress to rank budget $r_{\max}$ to obtain $\theta^{(g,t+1)}$; otherwise set $\theta^{(g,t+1)}\leftarrow\theta^{(g)}$.
@@ 4.1 Experimental Setup and Baselines
Item Setting
Backbone LLaMA-7B
Adapter LoRA
Tasks GLUE (RTE, QQP, MNLI, SST-2), MMLU
Clients ($K$) 8
Rounds ($T$) 30
Rank (hetero.) $\{r_{k}\}=[4,4,8,8,8,8,16,16]$
Training Optimizer: Adam; Learning rate $2\times 10^{-4}$
Local update steps per round: 10; Batch size: 128
Hardware Intel Xeon Gold 5318Y
[…]
Software Ubuntu 22.04.5; Python 3.10; CUDA 12.4
Table 1: Experimental configuration and system environment.
Hetero (Rank –) Homo
Task Padding FLoRA Ours $\boldsymbol{\Delta}$ Rank = 4 Rank = 8 Rank = 16
FedAvg Ours $\boldsymbol{\Delta}$ FedAvg Ours $\boldsymbol{\Delta}$ FedAvg Ours $\boldsymbol{\Delta}$
QNLI 95.81 96.10 96.96 +0.86 95.81 97.78 +1.97 96.50 98.47 +1.97 97.00 98.65 +1.65
RTE 95.45 97.20 99.71 +2.51 85.94 88.89 +2.95 85.62 90.05 +4.43 86.10 90.80 +4.70
QQP 86.72 87.10 87.97 +0.87 84.40 87.60 +3.20 83.10 85.30 +2.20 86.50 89.87 +3.37
MNLI(matched) 65.48 69.30 72.19 +2.89 70.20 73.05 +2.85 69.85 73.60 +3.75 70.10 74.00 +3.90
MNLI(mismatch) 64.90 68.80 71.62 +2.82 69.55 72.48 +2.93 69.20 73.05 +3.85 69.50 73.70 +4.20
SST-2 81.10 81.70 82.30 +0.60 82.06 82.14 +0.08 81.90 81.88 0.02-0.02 82.00 82.20 +0.20
MMLU 41.55 44.10 45.60 +1.50 44.00 46.30 +2.30 44.80 47.00 +2.20 45.25 47.60 +2.35
Avg. 75.86 77.47 79.48 +1.72 75.99 80.00 +2.32 75.00 78.94 +2.63 76.18 81.55 +2.91
Table 2: Federated aggregation results under rank-heterogeneous and rank-homogeneous clients. $\Delta$ denotes absolute improvement (percentage points) of Ours over the strongest baseline in each setting: $\max(\text{FLoRA},\text{Padding})$ for Hetero, and FedAvg for Homo (Rank=4/8/16).
@@
We evaluate federated parameter-efficient fine-tuning on GLUE-style and MMLU benchmarks Wang et al. (2018) Hendrycks et al. . We report results on QNLI, RTE, QQP, MNLI, and SST-2, which cover natural language inference, paraphrase identification, and sentiment classification. We use the official train/dev splits and report dev-set performance.
@@
We adopt LLaMA-7B Touvron et al. (2023) as the frozen backbone and insert LoRA adapters Hu et al. (2022) into linear projection layers. Only LoRA parameters are optimized while the backbone weights remain fixed.
@@
We simulate a cross-device FL setting with $K=8$ clients and run $T=30$ communication rounds. Unless otherwise stated, all clients participate in each round and the server aggregates client updates with weights proportional to local data sizes, i.e., $\alpha_{k}=n_{k}/\sum_{j}n_{j}$. To model structural heterogeneity induced by rank mismatch, we set client ranks as $\{r_{k}\}=[4,4,8,8,8,8,16,16]$. When a homogeneous setting is required, we use a shared rank $r_{k}=r$ for all clients. In an additional scalability study, we vary the number of participating clients $K$ while keeping the remaining hyperparameters consistent.
@@
Each client uses Adam with a fixed learning rate of $2\times 10^{-4}$ and batch size $128$. Within each communication round, clients perform a fixed number of local update steps ($10$) on mini-batches before uploading their LoRA updates. All remaining optimizer hyperparameters follow the default settings in our implementation. These hyperparameters are kept identical across methods unless the experiment explicitly controls a specific factor, ensuring fair comparisons. [^1] The above hyperparameters follow the experimental configuration summarized in Table 1.
@@
We compare the proposed aggregation strategy against representative alternatives under identical training budgets: (i) Zero-padding aligns heterogeneous LoRA updates by padding lower-rank factors to the maximum rank before weighted averaging; (ii) FedAvg performs standard weighted averaging of LoRA parameters in the homogeneous-rank setting; (iii) Stacking(FLoRA) concatenates client low-rank updates along the rank dimension to preserve the union of client subspaces.
@@ 4.2 Results under Rank and Data Heterogeneity
Rank heterogeneous clients.
Table 2 reports results under mismatched client ranks. Overall, Ours consistently outperforms zero-padding and FLoRA across all tasks. For example, it improves QNLI from 95.81% to 96.96% and QQP from 86.72% to 87.97%, and achieves notably larger gains on RTE and MNLI, which are more sensitive to cross-client directional mismatch.
These improvements stem from how different methods handle structural mismatch under rank heterogeneity. Zero-padding dilutes informative low-rank updates by enforcing a common high-rank parameterization, while FLoRA aggregates adapters in a monolithic manner. By contrast, Ours performs selective aggregation only on the shared component, leading to more stable and reliable updates under mismatched ranks.
Homogeneous rank sanity check.
To verify that the benefit of our method is not merely due to resolving dimensional incompatibility, we further compare it with FedAvg-style averaging when all clients use the same rank (Table 2). Even in this homogeneous setting, stacking remains consistently superior or comparable.
For rank $r=4$, Ours improves QQP by 3.20%, QNLI by 1.97%, and RTE by 2.95%, while maintaining comparable performance on SST-2. For rank $r=8$, Ours again improves QQP (2.20%), QNLI (1.97%), and RTE (4.43%), with only a negligible difference on SST-2 (-0.02%). For rank $r=16$, Ours yields further gains on QQP (3.37%), QNLI (1.65%), and RTE (4.70%), while slightly improving performance on SST-2 (0.20%). These results indicate that Ours acts as a subspace-preserving aggregator that remains beneficial even when the rank is fixed, rather than being a mere workaround for heterogeneous dimensions.
@@
We study the impact of the federated population size by varying the number of participating clients $K$ while keeping all other settings fixed. As shown in Figure 2, increasing $K$ generally improves test accuracy, reflecting the benefit of broader data coverage in larger federations. With very few clients (K=4K{=}4), aggregation is less effective due to limited diversity of update directions.
The performance trend is not strictly monotonic, and a temporary drop is observed at intermediate $K$ values (e.g., around K=16K{=}16), which is consistent with non-IID federated optimization. Comparing the two rank settings, lower ranks exhibit better robustness in the small-to-medium regime, whereas higher ranks achieve a higher accuracy ceiling as the client population grows. Overall, these results indicate that our method scales well with the federation size while exhibiting a non-trivial interaction between rank and population size.
@@
We further examine robustness under Non-IID data partitions controlled by a Dirichlet parameter $\alpha$ using the client-level accuracy standard deviation in Figure 3. Lower variance indicates more consistent performance across heterogeneous clients.
@@
At $\alpha=1.0$, stacking remains slightly more stable on RTE (0.088 vs. 0.092), but exhibits higher variance on SST-2 and QNLI. This suggests that when data heterogeneity is moderate, preserving a broader set of client-specific low-rank directions may amplify inter-client discrepancies on certain tasks, even though the shared model remains competitive. Importantly, when the partition approaches near-IID ($\alpha=10$), stacking again achieves uniformly lower variance across all tasks, showing that stacking does not inherently harm stability.
Figure 3: Client-level performance comparison between single-module and dual-model adapters under the same training budget. The decoupled design consistently improves accuracy across clients, validating the effectiveness of separating shared and private low-rank adaptations.
@@ 4.3 Privacy and Rank Sensitivity
Privacy and utility trade-off.
We apply DP-SGD during private training and report task accuracies under different privacy budgets $\epsilon$ (Figure 4). As $\epsilon$ increases (weaker privacy), the required noise scale decreases and performance improves.
Figure 4: Accuracy under different privacy budgets $\epsilon$. Fixed denotes applying structural constraints (selective stacking and rank control) on the shared module, while Unfixed corresponds to unconstrained aggregation over all updates. The fixed design exhibits substantially better stability under strong privacy constraints.
Figure 4 further shows that enforcing structural constraints on the shared module significantly improves robustness under strong privacy constraints. When $\epsilon$ is small, unconstrained aggregation suffers from severe performance degradation, whereas the fixed shared-module design maintains substantially higher accuracy.
@@
Table 3 analyzes how the DP noise affects performance. Without DP, smaller ranks can be more stable on some tasks, while overly large ranks may introduce redundancy and hurt aggregation. With DP ($\epsilon=1$), the optimal rank can shift due to noise, but small-to-medium ranks often remain a good trade-off.
@@
4.4 Evidence for Decoupled Benefit
Finally, we include direct evidence that introducing a dual-adapter (decoupled) design improves federated fine-tuning. Under matched training budgets, the dual-adapter achieves a higher average performance than a single-adapter baseline, indicating that dual modules better exploit complementary low-rank subspaces.
@@ 5 Conclusion
We studied federated fine-tuning of LLMs with LoRA under practical client heterogeneity, including differences in data distributions, system constraints, adapter ranks, and privacy requirements. While stacking-based aggregation provides a natural way to reconcile rank mismatch, we argue that stacking alone is insufficient, as it implicitly forces shared alignment on all client updates, suppressing personalization and becoming fragile under differential privacy noise.
We proposed the SDFLoRA framework, which structurally decomposes each client LoRA adapter into shared and private components. Building on this decomposition, we introduce selective stacking that applies stacking-based alignment only to the shared module while keeping private modules private and unaggregated. This configuration supports privacy-aware optimization by injecting DP noise exclusively into the shared module. Experiments on multiple benchmarks show that our approach outperforms representative federated LoRA baselines and achieves a better balance between utility and privacy under DP noise. In future work, we plan to extend this structural separation to other adapter families and investigate adaptive mechanisms for allocating shared versus private capacity based on client characteristics and task demands.